OpenStudy (anonymous):

Of 8 household chores, in how many ways can you do three-fourths of them?

OpenStudy (experimentx):

8C6

OpenStudy (anonymous):

or for someone who hasnt encountered binomials - "three fourths" of eight is 6 so we are choosing 6 chores, the first choice we have 8 options , then we have 7, etc etc so (8*7*6*5*4*3) , but then we have counted a bunch of choices more than once,( eg doing the washing up and then the hoovering is the same choice as the hoovering then the washing up_ so we have to divide that number by the different ways to arrange the 6 we chose: (8*7*6*5*4*3)/(6*5*4*3*2*1) which is, as experimentX says, 8C6 , which is 28

OpenStudy (anonymous):

in general, choosing r objects from a set of n objects there are nCr different possible combinations : nCr = (n!)/(r!(n-r)!) where the exclamation mark means "factorial" ( basically x! = x*(x-1)*(x-2)*....2*1 , for example 10! = 10*9*8*7*6*5*4*3*2*1 )

OpenStudy (anonymous):

N.B 0! = 1

OpenStudy (experimentx):

nice job explaining :D (no choosing 6 places out of 8)/(no of choosing 6 places)